# SD++: Enhancing Standard Definition Maps by Incorporating Road Knowledge using LLMs

### v1

@@ Abstract
High-definition maps (HD maps) are detailed and informative maps capturing lane centerlines and road elements. Although very useful for autonomous driving, HD maps are costly to build and maintain. Furthermore, access to these high-quality maps is usually limited to the firms that build them. On the other hand, standard definition (SD) maps provide road centerlines with an accuracy of a few meters. In this paper, we explore the possibility of enhancing SD maps by incorporating information from road manuals using LLMs. We develop SD++, an end-to-end pipeline to enhance SD maps with location-dependent road information obtained from a road manual. We suggest and compare several ways of using LLMs for such a task. Furthermore, we show the generalization ability of SD++ by showing results from both California and Japan.
@@ I Introduction
High-Definition (HD) maps play a critical role in autonomous driving systems, supporting key components such as localization, prediction, and planning. Their popularity stems from their precision in capturing road geometry and providing lane-level details, offering valuable context for a wide range of driving tasks. HD maps leverage advances in sensor fusion technology, incorporating multiple modalities that complement each other. For instance, 3D point cloud data from LiDAR is often enriched with semantic information derived from camera inputs, enabling a comprehensive understanding of lanes, junctions, and road segments.
@@
Despite their importance, the creation and maintenance of HD maps remain resource-intensive and challenging. Producing these maps requires extensive data collection using vehicles equipped with high-end sensors like LiDAR and RGB cameras. This raw data then undergoes sophisticated processing to extract detailed semantic features like lane boundaries. However, the process often involves significant manual annotation, which is both time-consuming and labor-intensive. Furthermore, HD maps require frequent updates to accommodate changes in road conditions, such as construction or layout modifications, making them costly and difficult to scale. Prior studies [1] highlight these operational challenges, emphasizing the need for more efficient and scalable solutions. To address these limitations, researchers have explored alternative approaches, such as generating HD maps directly from onboard sensor data [2, 3, 4, 5]. While promising, these methods often face challenges with generalization, limited range, and still require significant data collection.
@@
Figure 1: Overview of our proposed pipeline, which explores the potential of leveraging open-source map data, large language models, and official documents (e.g., highway design manuals) for mapping.
@@
Standard-Definition (SD) maps offer a cost-effective and scalable alternative. Open-source options like OpenStreetMap (OSM) [6] represent road networks using basic geographical coordinates, making them easy to scale. However, SD maps lack the fine-grained detail and accuracy [7] required for many autonomous driving applications, limiting their direct usability. Recent research [8, 9] has attempted to generate HD maps by integrating them with real-time sensor data, improving their usability for lane topology reasoning. Nevertheless, these methods still face challenges related to generalization and data dependency. Given that most existing approaches extract lane-level details primarily from visual cues, we explore whether SD maps can be enhanced using other publicly available resources, potentially narrowing the gap between SD and HD maps.
@@
In this study, we propose a novel approach to bridge the gap between HD and SD maps without relying on sensor data or manual annotation. Specifically, we explore the use of large language models (LLMs) to enhance SD maps using publicly available resources, such as highway design manuals [10] that provide detailed road design specifications. By leveraging these resources alongside OSM data, our method aims to generate enriched map representations that approximate the detail and utility of HD maps. Our key contributions are as follows:
@@
This work aims to provide a scalable and cost-effective alternative to traditional HD map creation while retaining its critical functionality for autonomous driving systems.
@@ III-B Road Manuals
Highway Design Manual (HDM) [10] is a comprehensive document that provides standardized guidelines and specifications for the design, construction, and maintenance of roadways and related infrastructure. It provides detailed parameters such as lane widths, shoulder dimensions, road alignments, and traffic control elements like signage and markings. These manuals are created and maintained by government transportation departments, such as the U.S. Department of Transportation (DOT) or state-level agencies to ensure safety, efficiency, and consistency in highway design across different regions.
@@
An example illustration from HDM is shown in Fig. 2
Figure 2: An example illustration from the Highway Design Manual (HDM)
@@ III-C Retrieval-Augmented Generation (RAG)
RAG is an approach in natural language processing that combines the retrieval of relevant information from an external knowledge base with the generation of text using a language model. Large documents are first broken down into smaller chunks. These chunks of text are converted to embeddings using an embedding model and stored in a vector database. Then, the embedding of a user query is compared with all the stored embeddings using similarity measures like the cosine similarity, and the document chunks corresponding to $k$, where most similar embeddings are retrieved. The retrieved documents are then appended to the user query and passed to the LLM to generate the final answer. This method enables the language model to access domain-specific information, improving its accuracy and relevance for specific tasks. RAG is beneficial for SD map enhancement due to the following reasons.
  • Incorporating Domain Knowledge: RAG allows the integration of external resources like highway design manuals or traffic regulations. By retrieving specific details, such as lane dimensions or intersection configurations, and providing them as context to the language model, RAG ensures the outputs adhere to real-world standards.
  • Enhancing Consistency: By retrieving consistent and structured information from SD maps or related documents, RAG can help create enhanced HD maps that follow established rules and guidelines, reducing errors or inconsistencies.
  • Dynamic Updates: As road conditions or standards change, RAG can retrieve the most up-to-date information, enabling HD map generation processes to adapt without retraining the language model.
  • Reducing Model Limitations: LLMs alone may struggle with domain-specific or granular tasks like accurate map enhancement. RAG bridges this gap by grounding the model’s outputs in retrieved, authoritative information.
@@ IV Methodology
In this section, we introduce two pipelines for generating enhanced Standard Definition (SD) maps: Direct Generation and Knowledge-Based Algorithmic Generation. The comparison of these methods and their results are presented in Section V. Inspired by the recent advancements in Large Language Models (LLMs) for text generation [23, 24], we explore their potential to enhance SD maps, given that SD maps, such as OpenStreetMap (OSM) data, are represented in text form. The Direct Generation method uses a straightforward approach by combining OSM input, LLM capabilities, and highway design manual (HDM) knowledge. In contrast, Knowledge-Based Algorithmic Generation incorporates Retrieval-Augmented Generation (RAG) techniques and additional post-processing steps, leveraging LLMs for specific tasks while addressing the limitations of direct text-based map generation.
@@ IV-A1 Process
  • Incorporating Highway Design Manual (HDM): The HDM is uploaded with instructions for the LLM to extract key road-related details such as lane width, shoulder widths, and other essential parameters for SD map enhancement.
@@ IV-B2 Incremental Generation (IG)
IV-B3 Autoregressive Generation (Auto)
Just like IG variant, Auto generates lanes one-by-one. But here, the pipeline is run in an autoregressive manner. That is, once we generate information for one lane, it is appended to the context for generating the next lane. Hence, the generated lanes benefit from the extracted information of other lanes.
In terms of cost, Auto $>$ IG $\gg$ OSG. We compare these variants in section V to see if the added expense is worth the performance gain.
@@ V-A Implementational Details
For the generation component of the RAG pipeline, we primarily use GPT-4o as the large language model (LLM). We also experiment with Llama [24], and provide a comparison of their performance in the following subsections.
@@ V-B Direct generation vs Algorithmic Generation
GPT-4o ($P1$) OSG 83 $32.04\pm 36.12$ $11.68\pm 22.17$ $0.28\pm 0.27$
GPT-4o ($P2$) OSG 100 $2.85\pm 0.4$ $0.23\pm 0.11$ $0.77\pm 0.11$
GPT-4o IG 100 $2.87\pm 0.37$ $0.26\pm 0.16$ $0.78\pm 0.11$
GPT-4o Auto 100 $3.01\pm 0.51$ $0.44\pm 0.29$ $0.76\pm 0.11$
Llama ($P1$) OSG 0 N/A N/A $00.00\pm 0$ 0.00
Llama ($P2$) OSG 50 $3.99\pm 1.47$ $2.99\pm 2.28$ $0.10\pm 0.08$
Llama IG 100 $3.12\pm 0.31$ $0.41\pm 0.51$ $0.78\pm 0.11$
Llama Auto 100 $3.32\pm 1.18$ $0.58\pm 0.74$ $0.72\pm 0.21$
TABLE I: Quantitative results on Argoverse 2 comparing various variants of SD++. N/A indicates no matching lane for chamfer distance calculation.
@@ V-C1 Evaluation method
We compare the output of SD++ with Argoverse ground truth using two metrics: Averaged Chamfer distance [31], and Recall. To calculate the Chamfer distance-based metric, the corresponding road in SD++ map for every road in Argoverse is established by matching their OSM way ID. The correspondence between predicted and ground-truth lanes is determined by minimizing the Chamfer distance. Table I shows the average of all these minimum distances. A SD++ lane prediction is considered correct if its Chamfer Distance to a lane in Argoverse 2 is less than 5 meters, as this threshold roughly corresponds to the typical road width. Following that definition of a correct lane, the recall is calculated as
@@ V-C2 Quantitative results
The quantitative evaluation of our approach on Argoverse is provided 2 in Table I. Our experiments reveal several key insights regarding the performance of different language models and the impact of prompt design.
First, we compare the performance of two large language models (LLMs), GPT-4o and Llama 3.1. Our results show that GPT-4o consistently outperforms Llama across various methods. Notably, GPT-4o achieves strong results simply by using an improved prompt, whereas Llama benefits more from an incremental generation approach. This difference in performance may be attributed to the model size, as we use the 7-billion-parameter version of Llama, which is significantly smaller than GPT-4o.
@@ V-C3 Qualitative result
Overall, our results highlight the crucial role of both prompt engineering and model selection in improving SD map enhancement. By refining prompt strategies and optimizing incremental processing, we can achieve more reliable and detailed lane structures.
@@ V-D Results from Japan
VI Conclusion
In this work, SD++ presents a novel map representation without needing sensor data. By leveraging LLMs and RAG, we utilize official documents like Highway Design Manuals to enhance existing SD maps with lane-level details, providing valuable context for downstream tasks. SD++ demonstrates generalizability across OSM data from different states and other countries. However, it has limitations; SD maps’ accuracy is not improved since no additional sensor data is used. Also, the performance of the SD++ depends notably on the road manuals used. The outputs will suffer if the road manuals do not contain adequate information. Future work will focus on improving intersection handling and possibly exploring live official web pages, such as Caltrans project data [36], to achieve more accurate results and keep road users updated with the latest information. Overall, we see significant potential for LLMs in advancing mapping technologies, especially when used as a for generating strong prior for autonomous driving applications.

### v2

@@ Abstract
High-definition maps (HD maps) are detailed and informative maps capturing lane centerlines and road elements. Although very useful for autonomous driving, HD maps are costly to build and maintain. Furthermore, access to these high-quality maps is usually limited to the firms that build them. On the other hand, standard definition (SD) maps provide road centerlines with an accuracy of a few meters. In this paper, we explore the possibility of enhancing SD maps by incorporating information from road manuals using LLMs. We develop SD++, an end-to-end pipeline to enhance SD maps with location-dependent road information obtained from a road manual. Our pipeline requires no sensor data input and only relies on road manuals and SD maps. We experiment several ways of using LLMs for map enhancement. Furthermore, we demonstrate the generalization ability of SD++ by showing results from six states in the United States and Japan. Code is available at https://github.com/AutonomousVehicleLaboratory/SDplusplus
@@ I Introduction
High-Definition (HD) maps play a critical role in autonomous driving systems, supporting key components such as localization, prediction, and planning through precise lane-level detail. HD maps leverage advances in sensor fusion technology, incorporating multiple modalities that complement each other to provide rich road context.
@@
Despite their importance, the creation and maintenance of HD maps remain resource-intensive and challenging. Producing these maps requires extensive data collection using vehicles equipped with high-end sensors like LiDAR. This raw data then undergoes sophisticated processing to extract detailed semantic features like lane boundaries. However, the process often involves significant manual annotation, which is both time-consuming and labor-intensive. Furthermore, HD maps require frequent updates to accommodate changes such as construction or layout modifications, making them costly and difficult to scale. Prior studies [1] highlight these operational challenges, emphasizing the need for more efficient and scalable solutions. To address these limitations, researchers have explored alternative approaches, such as generating HD maps directly from onboard sensor data [2, 3, 4, 5]. While promising, these methods often face challenges with generalization, limited range, and still require significant data collection.
@@
Figure 1: Overview of our proposed pipeline, which explores the potential of leveraging OpenStreetMap, LLMs, and official documents for mapping.
@@
Standard-Definition (SD) maps offer a cost-effective and scalable alternative. OpenStreetMap (OSM) [6] represent road networks using basic geographical coordinates, making them easy to scale. However, SD maps lack the fine-grained detail and accuracy [7] required for many autonomous driving applications, limiting their direct usability. Recent work [8, 9] has attempted to generate HD maps by integrating them with sensor data, improving their usability for lane topology reasoning. Nevertheless, these methods still face challenges related to generalization and data dependency. Given that most existing approaches extract lane-level details from visual cues, we explore the possibility of improving SD maps using other public resources to help close the gap between SD and HD maps.
@@
In this study, we propose a novel approach to bridge the gap between HD and SD maps without relying on any sensor data or manual annotation. Specifically, we explore the use of large language models (LLMs) to enhance SD maps using publicly available resources, such as highway design manuals (HDM) [10] that provide detailed road design specifications. By leveraging these resources alongside OSM data, our method aims to generate enriched map representations that approximate the detail and utility of HD maps. Our key contributions are as follows:
@@
This work aims to provide a scalable, cost-effective, and informative map prior while retaining its critical functionality for autonomous driving systems.
@@ III-B Road Manuals
Highway Design Manual (HDM) [10] is a comprehensive document that provides standardized guidelines and specifications for the design, construction, and maintenance of roadways and related infrastructure. It provides detailed parameters such as lane widths, shoulder dimensions, road alignments, and traffic control elements like signage and markings, shown in Fig. 2. These manuals are created and maintained by government transportation departments, such as the U.S. Department of Transportation (DOT) or state-level agencies to ensure safety, efficiency, and consistency in highway design across different regions. The contained information is also regional, e.g. in the US we can find HDMs at national, state and city level.
@@
Figure 2: An example illustration from the HDM
@@ III-C Retrieval-Augmented Generation (RAG)
RAG combines external information retrieval with language model generation. Large documents are split into chunks, embedded via an embedding model, and stored in a vector database. At inference, a user query is embedded and compared with stored embeddings (e.g., using cosine similarity) to retrieve the top-$k$ most relevant chunks. These are appended to the query and fed into the language model to produce a response. This approach allows the model to incorporate domain-specific knowledge, improving accuracy and relevance.
RAG is well-suited for SD map enhancement for several reasons:
  • Domain Knowledge Integration: RAG can retrieve relevant details from sources like highway design manuals, helping ensure outputs reflect real-world standards.
  • Improved Consistency: By grounding outputs in structured, authoritative data, RAG reduces inconsistencies in map generation.
  • Adaptability: RAG enables access to updated information without retraining, supporting dynamic map updates.
  • Overcoming LLM Limitations: It enhances the LLM’s ability to handle fine-grained, domain-specific tasks by providing relevant context.
@@ IV Methodology
In this section, we introduce two pipelines for generating enhanced Standard Definition (SD) maps: Direct Generation and Knowledge-Based Algorithmic Generation. The comparison of these methods and their results are presented in Section V. Inspired by the recent advancements in Large Language Models (LLMs) for text generation [23, 24], we explore their potential to enhance SD maps, given that SD maps, such as OpenStreetMap (OSM) data, are represented in text form. The Direct Generation method uses a straightforward approach by combining OSM input, LLM capabilities, and HDM knowledge. In contrast, Knowledge-Based Algorithmic Generation incorporates Retrieval-Augmented Generation (RAG) techniques and additional post-processing steps, leveraging LLMs for specific tasks while addressing the limitations of direct text-based map generation.
@@ IV-A1 Process
  • Incorporating HDM: The HDM is uploaded with instructions for the LLM to extract key road-related details such as lane width, shoulder widths, and other essential parameters for SD map enhancement.
@@ IV-B2 Iterative Generation (IG)
IV-B3 Autoregressive Generation (IG+Context)
Just like IG variant, IG+Context generates lanes one-by-one. But here, the pipeline is run in an autoregressive manner. That is, once we generate information for one lane, it is appended to the context for generating the next lane. Hence, the generated lanes benefit from the extracted information of other lanes.
In terms of cost, IG+Context $>$ IG $\gg$ OSG. We compare these variants in section V to see if the added expense is worth the performance gain.
@@ V-A Implementational Details
For the generation component of the RAG pipeline, we compare the performances of two famous Large Language Models (LLMs) - GPT-4o and Llama. We select these two models because while the former is proprietary, the latter is open-sourced.
@@ V-B Direct generation vs Algorithmic Generation
Baseline N/A 100 $3.53\pm 2.37$ $0.08$ $0.73$
GPT-4o ($P1$) OSG 100 $24.73\pm 36.62$ $0.14$ $0.39$
GPT-4o ($P2$) OSG 100 $2.52\pm 2.45$ $0.1$ $0.8$
GPT-4o IG 100 $2.54\pm 2.45$ $0.22$ $0.8$
GPT-4o IG+Context 100 $2.91\pm 2.46$ $0.08$ $0.78$
Llama ($P1$) OSG 16 $4.94\pm 8.09$ $0.77$ $0.01$
Llama ($P2$) OSG 65 $2.65\pm 2.13$ $0.38$ $0.18$
Llama IG 100 $2.5\pm 2.45$ $0.1$ $0.81$
Llama IG+Context 100 $2.5\pm 2.45$ $0.1$ $0.81$
TABLE I: Quantitative results on Argoverse 2 comparing various variants of SD++.
@@ V-C1 Evaluation method
We compare the output of SD++ with Argoverse ground truth using two metrics: Averaged Chamfer distance [31], and Recall. The reason for using Recall instead of precision is SD++ include larger area than Argoverse ground truth. To calculate the Chamfer distance-based metric, the corresponding road in SD++ map for every road in Argoverse is established by matching their OSM way ID. The correspondence between predicted and ground-truth lanes is determined by minimizing the Chamfer distance. Table I shows the average of all these minimum distances. A SD++ lane prediction is considered correct if its Chamfer Distance to a lane in Argoverse 2 is less than 5 meters, as this threshold roughly corresponds to the typical road width. Following that definition of a correct lane, the recall is calculated as
@@ V-C2 Quantitative results
The quantitative evaluation of our approach on Argoverse is provided 2 in Table I. Our experiments reveal several key insights regarding the performance of different language models and the impact of prompt design. The baseline method, A/B Street, uses a hand-crafted, rule-based approach to augment OSM. Interestingly, it performs quite well, which we attribute to the strong domain knowledge embedded by its designer.
First, we compare the performance of two large language models (LLMs), GPT-4o and Llama 3.3. Our results show that GPT-4o consistently outperforms Llama across various methods. Notably, GPT-4o achieves strong results simply by using an improved prompt, whereas Llama benefits more from an iterative generation approach. This difference in performance may be attributed to the model size, as we use the 7-billion-parameter version of Llama, which is significantly smaller than GPT-4o.
@@ V-C3 Qualitative result
Overall, our results highlight the crucial role of both prompt engineering and model selection in improving SD map enhancement. By refining prompt strategies and optimizing iterative processing, we can achieve more reliable and detailed lane structures.
@@ V-D Results from Japan
VI Limitation and Future Work
While our method demonstrates strong performance and generalization, it has several limitations. First, the accuracy of SD maps is not improved, as no sensor data is used. Second, the performance of SD++ depends heavily on the quality and completeness of the road manuals. If the manuals lack necessary details, such as specific lane dimensions or configurations, the output may be limited. In cases where OSM does not explicitly show certain road features—such as curved segments or uncommon layouts—SD++ cannot recover that missing information. Additionally, the reliability of our approach is influenced by the LLM’s output, which can occasionally be inconsistent or incorrect.
For future work, we plan to improve intersection handling and explore the use of live official sources, such as Caltrans project data [36], to improve accuracy and ensure the maps reflect recent changes. We also aim to extend SD++ as a prior for downstream tasks like trajectory prediction and online HD mapping.
VII Conclusion and Future Work
In this work, SD++ presents a novel map representation without needing sensor data. By leveraging LLMs and RAG, we utilize official documents to enhance existing SD maps with lane-level details, providing valuable context for downstream tasks. SD++ demonstrates generalizability across OSM data from different states and other countries. Overall, we see significant potential for LLMs in advancing mapping technologies, especially when used as a for generating strong prior for autonomous driving applications.